ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

16. Applications

1991 MS Classification: 76, 10 Vasiliev, Mikhail, Institute of Applied Mathematics, Moscow, Russia Automodelling solutions to the equations of the viscous heat conductive gas flows Truncated systems of the equations of viscous heat conductive gas flows are determined, and the flow on a semi-infinite flat plate is studied [1], using the general approach proposed by A. D. Bruno [2]. This approach is based upon the power exponent geometry including the Newton polyhedra. Here, for the first time it is applied to the system of PDE. References: 1. Vasiliev M. M. About the obtaining automodelling solutions to the equations of the viscous heat conductive gas flows / Institute of Appl. Math., Preprint 95, Moscow, 1997, (Rus). 2. Bruno A. D. Power exponent geometry in algebraic and differential equations. M., Nauka, (1998), (Rus).